There are 5 brothers having totai age of 50 years & 3 years difference between each brother so what is age of all 5 brothers individually?
Answer by Cromlix(4381)

Brothers' ages:
x, x+3, x+6, x+9, x+12
x+x+3+x+6+x+9+x+12 = 50
5x + 30 = 50
5x = 50 - 30
5x = 20
x = 4.
Brothers' ages
4yrs, 7yrs, 10yrs, 13yrs and 16yrs.
Hope this helps.
:-)